Kevin and tim have 56 coins in a jar. Some quarters and some nickels. It totals 5.80. How many are quarters and how many are nickels
Answer by oberobic(2304)

n = number of nickels
5n = value of the nickels in cents
q = number of quarters
25q = value of the quarters in cents
.
n + q = 56
n = 56 - q
.
5n + 25q = 580 cents
.
5*(56-q) +25q = 580
.
280 -5q +25q = 580
.
20q = 300
.
q = 15
.
n = 56 -q
n = 56 -15
n = 41
.
Check to see the vaue.
5(41) = 205
25(15) = 375
205 + 375 = 580
Correct.
.
Answer: They have 15 quarters and 41 nickels.
.
Done.